# Magnetic Polarisability of Octet Baryons via Lattice QCD

### v2

@@ V.1 Gauge ensembles
Table 3: Details of the PACS-CS ensembles used in this work. The lattice spacing of each ensemble is set using the Sommer scale with $r_{0}=0.49\,$fm. In all cases $\kappa_{s}^{\rm sea}=0.13640$ and $\kappa_{s}^{\rm val}=0.13665$ Menadue (2018). $N_{\rm con}$ describes the number of configurations.
@@
The four gauge ensembles used in this work are the four heaviest $2+1$-flavour dynamical gauge configurations provided by the PACS-CS collaboration Aoki et al. (2009) through the International Lattice Data Grid (ILDG) Di Renzo (2024). The configurations have a range of degenerate up and down quark masses while the strange quark mass is fixed. The strange quark mass of the ensembles which corresponds to $\kappa_{s}=0.13640$ does not extrapolate to the physical kaon mass Menadue et al. (2012). Use of $\kappa_{s}=0.13665$ for the valence strange quark mass produces the correct value for the kaon mass extrapolated to the physical point Menadue (2018). We note that the mass of the strange quarks in the sea remain at the heavier mass.
Each of the ensembles are a $32^{3}\times 64$ lattice. The gauge action is the Iwasaki gauge action and the clover fermion action with $C_{\rm SW}=1.715$ is the background field corrected clover fermion action Bignell et al. (2019) which is tuned to remove the non-physical magnetic-field induced additive mass renormalisation. Details of the ensembles are summarised in Tab. 3.
As only the valence quarks interact with the background magnetic field the ensembles are electro-quenched. While it is possible to include the background field in the process of generating each gauge field configuration Fiebig et al. (1989), this would require a separate Monte-Carlo simulation at each field strength. Such separate simulations would remove the correlated QCD fluctuations which are otherwise efficiently removed through the ratio in Eq. (51). Without this correlation, a very significant increase in statistics would be required. Instead we preserve the correlations and estimate the small corrections for electro-quenching through chiral effective field theory in the process of extrapolation to the physical point.
@@ V.2 Baryon interpolating fields
where $C$ is the charge-conjugation matrix. The interpolating fields of the other outer octet baryons may be easily obtained through appropriate substitution of doubly and singly represented quark flavours. This is the form of the interpolating field used throughout this work.
Such interpolating fields, utilised purely with traditional gauge-covariant Gaussian smearing are ineffective at isolating the baryon ground state in a uniform background field Primer et al. (2014); Deshmukh and Tiburzi (2018); Bignell et al. (2020); Bruckmann et al. (2017); Chang et al. (2015). The uniform background field breaks the spatial symmetry and Landau-mode physics presents at both the quark and hadronic levels, and this must be accommodated.
@@ V.3 Quark operators
Asymmetric source and sink operators have been shown to improve the overlap of the lowest energy eigenstates of baryons in a magnetic field Bignell et al. (2020). Following this work we utilise standard Gaussian smearing at the source and a low-lying eigenmode projection at the sink. The low-mode sink projection employs QED+QCD eigenmodes such that the quark propagators are sensitive to the dynamics of nontrivial electric charges in the magnetic field. As the QCD+QED eigenmodes contain the effects of the breaking of spatial symmetry by the uniform field in the $z$-direction, these modes further aid in the isolation of the lowest-energy baryon in the correlation functions. Finally, we include a U(1) projection of the colour-singlet baryon state for charged baryons rather than a traditional Fourier projection. This ensures the isolation of the lowest-Landau-level. Each of these steps is described in detail below.
@@ V.3.1 Link smearing
In constructing the smeared source and sink projections described below, stout link smearing Morningstar and Peardon (2004) is utilised on the spatially-oriented gauge links. 10 smearing sweeps are applied with an isotropic smearing parameter of $\alpha_{stout}=0.1$. These gauge links are used in the process of delta-function source smearing and in the calculation of the sink projection via low-lying eigenmodes of the lattice Laplacian.
@@ V.3.2 Quark propagator source
At all quark masses, $\alpha=0.7$ is used for the Gaussian smearing. The number of gauge-invariant Gaussian smearing sweeps considered is quark mass dependent, with smaller numbers of sweeps associated with heavier quark masses. In tuning the smearing to optimise the onset of early effective-mass plateaus Bignell et al. (2020), 150 through 350 sweeps are utilised.
@@ V.3.4 Quark propagator sink
where $U_{\mu}(\vec{x})$ are the full SU(3)$\times$×U(1) gauge links described in Sec. III. Again, stout-links are used in constructing the eigenmodes.
Due to the two dimensional nature of the Laplacian, the modes are calculated on each $(z,t)$-slice of the lattice independently. Considering the four-dimensional coordinate space representation of the eigenmode with $\vec{r}=(x,y,z)$
@@
we form a coordinate-space projection operator
@@
which may then be applied to the quark propagator at the sink. As we are looking for low-energy states, we utilise only the $n$ lowest-lying eigenmodes of the Laplacian. It is shown in Ref. Bignell et al. (2020), that including too few modes results in a noisy hadron correlation function in much the same manner as applying too many sweeps of traditional sink smearing. As such, the number of modes is chosen to be large enough to minimise the noise of the correlation function, but small enough to retain the focus on the aforementioned low-energy physics. The work of Ref. Bignell et al. (2020) found that $n=96$ modes provides balance to these two effects and is what we use here.
@@ V.4 Hadronic projection
Due to the colour singlet nature of the baryon, we need only project the eigenmodes of the U(1) Laplacian rather than the full lattice Laplacian used for the sink. Again we follow the formalism of Ref. Bignell et al. (2020).
@@
In the infinite-volume continuum limit, the lowest Landau mode has a Gaussian form, $\psi_{\vec{B}}\sim e^{-\absolutevalue{q_{B}e\,B}\,\left(x^{2}+y^{2}\right)/4}$. However, in a finite volume the periodicity of the lattice causes the wave function’s form to be altered Tiburzi and Vayl (2013); Bignell et al. (2020). As such, we instead calculate the lattice Landau eigenmodes using the two-dimensional U(1) lattice Laplacian in an analogous way to Eq. (54) Bignell et al. (2018). Here $U_{\mu}$ contains only the $U(1)$ phases appropriate to the background magnetic field quantised on the lattice.
@@
In evaluating Eq. (60), one can also consider the case of fixing $n=1$, and considering only one of the degenerate eigenmodes through an optimisation procedure. To ensure excellent overlap with the source, one rotates the $U(1)$ eigenmode basis to maximise overlap with the baryon source. Recalling the baryon delta-function source in the $x$-$y$ plane $\rho(x,y)=\delta_{x0}\,\delta_{y0}$, one proceeds to optimise the overlap of the source with the first mode $i=1$ by maximising the value of $\left|\innerproduct{\rho\,}{\,\psi_{i=1,\vec{B}}}\right|^{2}$. An optional phase is then applied such that $\psi_{i=1,\vec{B}}(0,0)$ is purely real at the source point. Illustrations of these eigenmodes are provided in Ref. Bignell et al. (2020).
Ref. Bignell et al. (2020) found that results using the one unrotated degenerate mode where the Landau wave function’s peak is centred on the quark source were almost indistinguishable to those using an optimised linear combination of the degenerate modes. As such, we perform the final-state charged baryon projection using the single degenerate mode peaked at the source. For large field strengths, the probability density of the projection mode has a Gaussian shape. As the field strength decreases the width of the Gaussian increases, making the probability distribution flatter as zero field strength is approached. Landau mode probability densities for charge $|q_{B}|=1$ baryons in fields strengths $|q_{d}|=1$ and $2$ are illustrated in Fig. 2.
Figure 2: Landau mode probability densities for charge $|q_{B}|=1$ baryons at our two lowest field strengths. Recalling $n=|k_{d}\,{q_{B}}/{q_{d}}|$, the $n=3$ (left) and $n=6$ (right) probability densities are illustrated for field strengths $|q_{d}|=1$ and $2$ respectively. The degeneracy of each mode is $n$; the mode best aligned with the fermion source is selected.
@@ V.5 Statistics
As periodic boundary conditions are used in all four dimensions for the gauge field generation, one can exploit the associated translational invariance of the gauge fields. A quark source can be placed at any position on the lattice and then circularly cycled to the standard source position of $(x,y,z,t)=(1,1,1,16)$. This enables additional sampling of the full gauge field.
Further, the two dimensional nature of the lattice Laplacian operator allows the eigenmodes for the sink projection to be re-used when the gauge field is cycled solely in the time direction. Hence, we increase our statistics on the PACS-CS ensembles by considering four random spatial sources at $t=16$. The gauge field is then circularly cycled in the temporal direction by an eighth of the lattice time extent (eight slices in our case) for each random source. This results in a further increase in statistics by a factor of eight. Together random sources and time-direction cycles increase our statistics by a factor of 32. These multiple samples are binned and averaged as a single configuration estimate in the error analysis.
@@ VI Fitting
As discussed in Sec. IV, we construct the spin-field aligned and anti-aligned correlation functions
@@
which aggregates the positive and negative field strengths together to remove the magnetic moment term from the energy expansion of Eq. (47). Due to this aggregation, any reference to the magnetic field strength from this point refers to the aggregated positive field strength.
Recalling Eq. IV, we then construct the magnetic polarisability energy shift
@@
Fitting to the above ratio in the long time limit and repeating this as a function of our two positive field strengths provides access to the magnetic polarisability.
In practice it is simple to fit in terms of the field strength quanta $k_{d}$. Using the quantisation condition of Eq. (46)
@@
we substitute for $e\,B$ in Eq. VI to re-write $\delta E_{\beta}$ in terms of the field strength quanta
@@
$=\frac{2\pi}{N_{x}\,N_{y}\,a^{2}}\frac{1}{2\,m}\absolutevalue{\frac{q_{B}}{q_{d}}\,k_{d}}-\frac{1}{2\,\alpha}\left[\frac{2\pi}{N_{x}\,N_{y}\,a^{2}}\right]^{2}\frac{1}{q_{d}^{2}}\,\beta\,k_{d}^{2},$ ((67))
@@
It is important to note the distinction between $q_{B}$ which is the charge of the baryon in question and $q_{d}$, the down quark charge introduced in the magnetic field quantisation condition. Here $\alpha=\frac{e^{2}}{4\pi}$ is the fine structure constant. For convenient future reference, we have defined $L(k_{d},m)$ the Landau term and $C$ the remaining coefficient to the magnetic polarisability term.
Previous fitting procedures Bignell et al. (2018, 2020) for the magnetic polarisability in the background field method followed the following steps
  • Consider all field strengths under investigation and select a common fit window for the magnetic polarisability energy shift. Fit the energy shift to determine $\delta E_{\beta}(k_{d})$.
@@
  • Fit $\left[\delta E_{\beta}(k_{d})-L(k_{d},m)\right]/C$ as a function of the integer $k_{d}^{2}$ to obtain the magnetic polarisability.
Specifically, fits for $\delta E_{\beta}(k_{d})$ were judged by the full covariance matrix $\chi^{2}_{\rm dof}$ coefficient, estimated by the jackknife method Efron (1979). We required $\chi^{2}_{\rm dof}<1.2$ at both field strengths with the start and end points of the fit window being common for each field strength. These criteria are designed to maximise the cancellation of correlated QCD fluctuations.
Figure 3: A plateau fit for a proton at $m_{\pi}=296\,$MeV. This effective energy plateau fit to $\delta E_{\beta}(B,t)$ is an example of a fit that looks acceptable, but is in fact fit too early due to underlying excited state contamination.
@@ VI.1 Fitting issues
This method presents a couple of issues. It makes the assumption that once the $\delta E_{\beta}(k_{d},t)$ energy shift plateaus, the underlying correlation functions used to construct the ratio are also independently exhibiting plateau-like behaviour. That is an assumption that a plateau in $\delta E_{\beta}(k_{d},t)$ implies single state isolation of all underlying correlation functions. We have seen that this is not necessarily the case.
Consider the plateau plot for the proton in Fig. 3 at $\kappa=0.13770$, $m_{\pi}=296\,$MeV. The selected fit window [21, 27] appears very reasonable. The $\chi^{2}_{\rm dof}$ values are both acceptable, the energy shift appears plateau-like at each field strength and we don’t appear to be fitting significant noise.
However, we need to check that the underlying correlators that form the ratio $R(k_{d},t)$ of Eq. (51) have all reached single state isolation.
@@
which reaches a linear form in the long time limit once single-state isolation has been reached. As such, the $\chi^{2}_{\rm dof}$ of a linear fit to $\log G$ is an excellent metric for single state isolation of the underlying correlator.
@@
To provide an easy interface to the vast number of $\chi^{2}_{\rm dof}$ values which need to be checked, we produce a series of heat maps. The heat maps corresponding to the underlying correlators of the proton magnetic polarisability are shown in Fig. 4.
@@
Each plot corresponds to a different underlying correlator. We must consider the zero-field correlator in addition to the aligned $G_{\uparrow\uparrow}$ and anti-aligned $G_{\uparrow\downarrow}$ correlators at each field strength. In each case we consider all windows which may be of interest. The heat maps clearly show that $t_{s}=21$, the starting time slice chosen in Fig. 3 is too early.
@@
Noting that the $\chi^{2}_{\rm dof}$ has a distribution, our absolute criteria for the choice of a starting time slice is the first three fit windows must satisfy the criteria $\chi^{2}_{\rm dof}<2.5$. We regard this criteria as a cut to ensure the fit windows considered in the averaging procedure described below are relevant and do not adversely affect the final result. At the same time, it is desirable to consider several meritorious fit windows such that an average weighted by the $\chi^{2}_{\rm dof}$ may be considered. The details of the weighted averaging are provided in subsection VI.3.
@@
The criteria $\chi^{2}_{\rm dof}<2.5$ corresponds to the three leftmost boxes in a row being blue, green, or yellow and this corresponds to plateau-like behaviour across 5 time slices. Earliest possible starting time slices in each case according to the criteria have been highlighted with green text in the column denoting $t_{s}$.
@@
Based on this, we see that we cannot begin fitting until time slice $t=24$, much later than $t=21$ which we selected based only on the plateau behaviour of the polarisability energy shift. This highlights the importance of checking the single state isolation of the underlying correlators. Pushing the fit later here changes the result significantly. We do note that in many cases, the polarisability energy shift does not plateau until the underlying correlators have reached single-state isolation, however it is important to check due to cases such as the proton presented here.
@@
Finally, we do not require the fit for $\log G$ to maintain a good quality of fit indefinitely as the correlation function can suffer a loss of signal. Instead, we use the measure as a metric for excited state contamination. We need only determine when $\log G$ becomes linear. As an example, consider the spin-field aligned, $k_{d}=2$, $\chi^{2}_{\rm dof}$ heat map for the neutron at $\kappa=0.13700$, $m_{\pi}=701\,$MeV shown in Fig. 5. Here we see two clear regions. The criteria is satisfied for fit windows $[t_{s},t_{f}]=[22,28],\ [23,28]$ and 24,28]24,28]24 , 28 ]. Then again at $t_{s}=28$ onwards. In this case, we take the earliest start point to be $t_{s}=22$, as the $\chi^{2}_{\rm dof}$ suggests clear single state isolation over 7 time slices. This scale is long enough to eliminate the possibility of a short-lived false plateau associated with an excited state. Further inspection reveals the source of the red and black blocks in the heat map commencing at $t_{f}=29$ is noise due to a loss of signal.
@@
Figure 5: $\chi^{2}_{\rm dof}$ heat map for fits to $\log G$ for all fit windows $t_{s}\to t_{f}$ of the spin-field aligned correlator at the second field strength. Here we see the behaviour exhibited in some cases in which the quality of $\log G$ fit is good in two places and unsatisfactory in between. This example is a neutron at $\kappa=0.13700$, $m_{\pi}=701$ MeV. Further inspection reveals the red quadrant is associated with a loss of signal in the correlator.

### v3

@@ V.1 Gauge ensembles
Table 3: Details of the PACS-CS ensembles used in this work. The lattice spacing of each ensemble is set using the Sommer scale with $r_{0}=0.4921(64)(+74)(-2)\,$fm. In all cases $\kappa_{s}^{\rm sea}=0.13640$ and $\kappa_{s}^{\rm val}=0.13665$ Menadue (2018). $N_{\rm con}$ describes the number of configurations.
@@
The four gauge ensembles used in this work are the four heaviest of five $2+1$-flavor dynamical gauge configurations provided by the PACS-CS collaboration Aoki et al. (2009) through the International Lattice Data Grid (ILDG) Di Renzo (2024). The configurations have a range of degenerate up and down quark masses while the strange quark mass is fixed.
The ensembles were generated in the absence of a background magnetic field. As a result, the sea quarks are blind to the magnetic field and the ensembles may be regarded as electroquenched. On the fifth and lightest ensemble, we encounter uncertainties which do not respond to increased statistics, hinting at an exceptional configuration problem associated with the electroquenching of the light sea-quark sector. For this reason, it has been omitted.
The strange-quark mass of the ensembles which corresponds to $\kappa_{s}=0.13640$ does not extrapolate to the physical kaon mass Menadue et al. (2012). Use of $\kappa_{s}=0.13665$ for the valence strange quark mass produces the correct value for the kaon mass extrapolated to the physical point Menadue (2018). We note that the mass of the strange quarks in the sea remain at the heavier mass.
Each of the ensembles considered are a $32^{3}\times 64$ lattice. The gauge action is the Iwasaki gauge action and the clover fermion action with $C_{\rm SW}=1.715$ is the background-field-corrected clover fermion action Bignell et al. (2019) which is tuned to remove the unphysical magnetic-field-induced additive mass renormalization. Details of the ensembles are summarized in Table 3.
As only the valence quarks interact with the background magnetic field the ensembles are electroquenched. While it is possible to include the background field in the process of generating each gauge-field configuration Fiebig et al. (1989), this would require a separate Monte Carlo simulation at each field strength. Such separate simulations would remove the correlated QCD fluctuations which are otherwise efficiently removed through the ratio in Eq. (51). Without this correlation, a very significant increase in statistics would be required. Instead we preserve the correlations and estimate the small corrections for electroquenching through chiral effective field theory in the process of extrapolation to the physical point.
The pion masses obtained by the PACS-CS Collaboration in their renormalisation scheme are used as a label for their ensembles. We denote these masses $m_{\pi}^{\rm PACS-CS}$ as in Table 3. Our analysis uses our pion masses calculated in the Sommer scheme Sommer (1994) with PACS-CS $r_{0}$ values from Table XII in Ref. Aoki et al. (2009). We also use their determination of the physical value $r_{0}=0.4921(64)(+74)(-2)\,$fm to set the lattice spacing for each ensemble. The pion masses obtained in this scheme are denoted $m_{\pi}$ and are given in Table 6.
@@ V.2 Baryon interpolating fields
where $C$ is the charge-conjugation matrix. The interpolating fields of the other outer octet baryons may be easily obtained through appropriate substitution of doubly and singly represented quark flavors. This is the form of the interpolating field used throughout this work.
Such interpolating fields, utilized purely with traditional gauge-covariant Gaussian smearing are ineffective at isolating the baryon ground state in a uniform background field Primer et al. (2014); Deshmukh and Tiburzi (2018); Bignell et al. (2020); Bruckmann et al. (2017); Chang et al. (2015). The uniform background field breaks the spatial symmetry and Landau-mode physics presents at both the quark and hadronic levels, and this must be accommodated.
@@ V.3 Quark operators
Asymmetric source and sink operators have been shown to improve the overlap of the lowest-energy eigenstates of baryons in a magnetic field Bignell et al. (2020). Following this work we utilize standard Gaussian smearing at the source and a low-lying eigenmode projection at the sink. The low-mode sink projection employs QED+QCD eigenmodes such that the quark propagators are sensitive to the dynamics of nontrivial electric charges in the magnetic field. As the QCD+QED eigenmodes contain the effects of the breaking of spatial symmetry by the uniform field in the $z$-direction, these modes further aid in the isolation of the lowest-energy baryon in the correlation functions. Finally, we include a U(1) projection of the color-singlet baryon state for charged baryons rather than a traditional Fourier projection. This ensures the isolation of the lowest Landau-level. Each of these steps is described in detail below.
@@ V.3.1 Link smearing
In constructing the smeared source and sink projections described below, stout link smearing Morningstar and Peardon (2004) is utilized on the spatially oriented gauge links. Ten smearing sweeps are applied with an isotropic smearing parameter of $\alpha_{stout}=0.1$. These gauge links are used in the process of $\delta$-function source smearing and in the calculation of the sink projection via low-lying eigenmodes of the lattice Laplacian.
@@ V.3.2 Quark propagator source
At all quark masses, $\alpha=0.7$ is used for the Gaussian smearing. The number of gauge-invariant Gaussian smearing sweeps considered is quark mass dependent, with smaller numbers of sweeps associated with heavier quark masses. In tuning the smearing to optimize the onset of early effective-mass plateaus Bignell et al. (2020), 150-350 sweeps are utilized.
@@ V.3.4 Quark propagator sink
where $U_{\mu}(\vec{x})$ are the full SU(3)$\times$×U(1) gauge links described in Sec. III. Again, stout links are used in constructing the eigenmodes.
Due to the two-dimensional nature of the Laplacian, the modes are calculated on each $(z,t)$-slice of the lattice independently. Considering the four-dimensional coordinate space representation of the eigenmode with $\vec{r}=(x,y,z)$
@@
we form a coordinate space projection operator
@@
which may then be applied to the quark propagator at the sink. As we are looking for low-energy states, we utilize only the $n$ lowest-lying eigenmodes of the Laplacian. It is shown in Ref. Bignell et al. (2020), that including too few modes results in a noisy hadron correlation function in much the same manner as applying too many sweeps of traditional sink smearing. As such, the number of modes is chosen to be large enough to minimize the noise of the correlation function, but small enough to retain the focus on the aforementioned low-energy physics. The work of Ref. Bignell et al. (2020) found that $n=96$ modes provides balance to these two effects and is what we use here.
@@ V.4 Hadronic projection
Due to the color-singlet nature of the baryon, we need only project the eigenmodes of the U(1) Laplacian rather than the full lattice Laplacian used for the sink. Again we follow the formalism of Ref. Bignell et al. (2020).
@@
In the infinite-volume continuum limit, the lowest Landau mode has a Gaussian form, $\psi_{\vec{B}}\sim e^{-\absolutevalue{q_{B}e\,B}\,\left(x^{2}+y^{2}\right)/4}$. However, in a finite volume the periodicity of the lattice causes the wave function’s form to be altered Tiburzi and Vayl (2013); Bignell et al. (2020). As such, we instead calculate the lattice Landau eigenmodes using the two-dimensional U(1) lattice Laplacian in an analogous way to Eq. (54) Bignell et al. (2018). Here $U_{\mu}$ contains only the U(1) phases appropriate to the background magnetic field quantized on the lattice.
@@
In evaluating Eq. (60), one can also consider the case of fixing $n=1$ and considering only one of the degenerate eigenmodes through an optimization procedure. To ensure excellent overlap with the source, we rotate the $U(1)$ eigenmode basis to maximize overlap with the baryon source. Recalling the baryon delta-function source in the $x$-$y$ plane $\rho(x,y)=\delta_{x0}\,\delta_{y0}$, we proceed to optimize the overlap of the source with the first mode $i=1$ by maximizing the value of $\left|\innerproduct{\rho\,}{\,\psi_{i=1,\vec{B}}}\right|^{2}$. An optional phase is then applied such that $\psi_{i=1,\vec{B}}(0,0)$ is purely real at the source point. Illustrations of these eigenmodes are provided in Ref. Bignell et al. (2020).
For large field strengths, the probability density of the projection mode has a Gaussian shape. As the field strength decreases the width of the Gaussian increases, making the probability distribution flatter as zero-field strength is approached. Landau-mode probability densities for charge $|q_{B}|=1$ baryons in fields strengths $|k_{d}|=1$ and $2$ are illustrated in Fig. 2.
Figure 2: Landau-mode probability densities for charge $|q_{B}|=1$ baryons at our two lowest field strengths. Recalling $n=|k_{d}\,{q_{B}}/{q_{d}}|$, the $n=3$ (left) and $n=6$ (right) probability densities are illustrated for field strengths $|k_{d}|=1$ and $2$ respectively. The degeneracy of each mode is $n$.
@@ V.5 Statistics
As periodic boundary conditions are used in all four dimensions for the gauge-field generation, one can exploit the associated translational invariance of the gauge fields. A quark source can be placed at any position on the lattice and then circularly cycled to the standard source position of $(x,y,z,t)=(1,1,1,16)$. This enables additional sampling of the full gauge field.
Further, the two-dimensional nature of the lattice Laplacian operator allows the eigenmodes for the sink projection to be reused when the gauge field is cycled solely in the time direction. Hence, we increase our statistics on the PACS-CS ensembles by considering four random spatial sources at $t=16$. The gauge field is then circularly cycled in the temporal direction by an eighth of the lattice time extent (eight slices in our case) for each random source. This results in a further increase in statistics by a factor of 8. Together random sources and time-direction cycles increase our statistics by a factor of 32. These multiple samples are binned and averaged as a single configuration estimate in the error analysis.
@@ VI Fitting
As discussed in Sec. IV, we construct the spin-field aligned and antialigned correlation functions
@@
which aggregates the positive and negative field strengths together to remove the magnetic moment term from the energy expansion of Eq. (47). Due to this aggregation, any reference to the magnetic-field strength from this point refers to the aggregated positive field strength.
Recalling Eq. IV, we then construct the magnetic polarizability energy shift
@@
Fitting to the above ratio in the long time limit and repeating this as a function of our two positive field strengths provides access to the magnetic polarizability.
In practice, it is simple to fit in terms of the field strength quanta $k_{d}$. Using the quantization condition of Eq. (46)
@@
we substitute for $e\,B$ in Eq. VI to rewrite $\delta E_{\beta}$ in terms of the field strength quanta
@@
$=\frac{2\pi}{N_{x}N_{y}a^{2}}\frac{1}{2m}\absolutevalue{\frac{q_{B}}{q_{d}}k_{d}}-\frac{1}{2\alpha}\left[\frac{2\pi}{N_{x}N_{y}a^{2}}\right]^{2}\frac{1}{q_{d}^{2}}\beta k_{d}^{2},$
@@
It is important to note the distinction between $q_{B}$ which is the charge of the baryon in question and $q_{d}$, the down quark charge introduced in the magnetic-field quantization condition. Here $\alpha=\frac{e^{2}}{4\pi}$ is the fine structure constant. For convenient future reference, we have defined $L(k_{d},m)$ the Landau term and $C$ the remaining coefficient to the magnetic polarizability term.
Previous fitting procedures Bignell et al. (2018, 2020) for the magnetic polarizability in the background-field method followed the following steps
  1. (i)
    Consider all field strengths under investigation and select a common fit window for the magnetic polarizability energy shift. Fit the energy shift to determine $\delta E_{\beta}(k_{d})$.
  2. (ii)
@@
  3. (iii)
    Fit $\left[\delta E_{\beta}(k_{d})-L(k_{d},m)\right]/C$ as a function of the integer $k_{d}^{2}$ to obtain the magnetic polarizability.
Specifically, fits for $\delta E_{\beta}(k_{d})$ were judged by the full covariance matrix $\chi^{2}_{\rm d.o.f.}$ coefficient, estimated by the jackknife method Efron (1979). We required $\chi^{2}_{\rm d.o.f.}<1.2$ at both field strengths with the start and end points of the fit window being common for each field strength. These criteria are designed to maximise the cancellation of correlated QCD fluctuations.
Figure 3: A plateau fit for a proton at $m_{\pi}=296\,$MeV. This effective energy plateau fit to $\delta E_{\beta}(B,t)$ is an example of a fit that looks acceptable, but is in fact fit too early due to underlying excited-state contamination.
@@ VI.1 Fitting issues
This method presents a couple of issues. It makes the assumption that once the $\delta E_{\beta}(k_{d},t)$ energy shift plateaus, the underlying correlation functions used to construct the ratio are also independently exhibiting plateaulike behavior. That is an assumption that a plateau in $\delta E_{\beta}(k_{d},t)$ implies single-state isolation of all underlying correlation functions. We have seen that this is not necessarily the case.
Consider the plateau plot for the proton in Fig. 3 at $\kappa=0.13770$, $m_{\pi}=296\,$MeV. The selected fit window [21, 27] appears very reasonable. The $\chi^{2}_{\rm d.o.f.}$ values are both acceptable, the energy shift appears plateaulike at each field strength, and we do not appear to be fitting significant noise.
However, we need to check that the underlying correlators that form the ratio $R(k_{d},t)$ of Eq. (51) have all reached single-state isolation.
@@
which reaches a linear form in the long time limit once single-state isolation has been reached. As such, the $\chi^{2}_{\rm d.o.f.}$ of a linear fit to $\log G$ is an excellent metric for single-state isolation of the underlying correlator.
@@
To provide an easy interface to the vast number of $\chi^{2}_{\rm d.o.f.}$ values which need to be checked, we produce a series of heat maps. The heat maps corresponding to the underlying correlators of the proton magnetic polarizability are shown in Fig. 4.
@@
Each plot corresponds to a different underlying correlator. We must consider the zero-field correlator in addition to the aligned $G_{\uparrow\uparrow}$ and antialigned $G_{\uparrow\downarrow}$ correlators at each field strength. In each case we consider all windows which may be of interest. The heat maps clearly show that $t_{s}=21$, the starting time slice chosen in Fig. 3, is too early.
@@
Noting that the $\chi^{2}_{\rm d.o.f.}$ has a distribution, our absolute criteria for the choice of a starting time slice is the first three fit windows must satisfy the criteria $\chi^{2}_{\rm d.o.f.}<2.5$. We regard this criteria as a cut to ensure the fit windows considered in the averaging procedure described below are relevant and do not adversely affect the final result. At the same time, it is desirable to consider several meritorious fit windows such that an average weighted by the $\chi^{2}_{\rm d.o.f.}$ may be considered. The details of the weighted averaging are provided in subsection VI.3.
@@
The criteria $\chi^{2}_{\rm d.o.f.}<2.5$ corresponds to the three leftmost boxes in a row being blue, green, or yellow and this corresponds to plateaulike behavior across five time slices. Earliest possible starting time slices in each case according to the criteria have been highlighted with green text in the column denoting $t_{s}$.
@@
Based on this, we see that we cannot begin fitting until time slice $t=24$, much later than $t=21$ which we selected based only on the plateau behavior of the polarizability energy shift. This highlights the importance of checking the single-state isolation of the underlying correlators. Pushing the fit later here changes the result significantly. We do note that, in many cases, the polarizability energy shift does not plateau until the underlying correlators have reached single-state isolation, however it is important to check due to cases such as the proton presented here.
@@
Finally, we do not require the fit for $\log G$ to maintain a good quality of fit indefinitely as the correlation function can suffer a loss of signal. Instead, we use the measure as a metric for excited state contamination. We need only determine when $\log G$ becomes linear. As an example, consider the spin-field aligned, $k_{d}=2$, $\chi^{2}_{\rm d.o.f.}$ heat map for the neutron at $\kappa=0.13700$, $m_{\pi}=701\,$MeV shown in Fig. 5. Here we see two clear regions. The criteria is satisfied at starting time slices $t_{s}=22,23,$ and $24$, then again at $t_{s}=28$ onward. In this case, we take the earliest start point to be $t_{s}=22$, as the $\chi^{2}_{\rm d.o.f.}$ suggests clear single-state isolation over seven time slices. This scale is long enough to eliminate the possibility of a short-lived false plateau associated with an excited state. Further inspection reveals the source of the red and black blocks in the heat map commencing at $t_{f}=29$ is noise due to a loss of signal.
@@
Figure 5: $\chi^{2}_{\rm d.o.f.}$ heat map for fits to $\log G$ for all fit windows $t_{s}\to t_{f}$ of the spin-field aligned correlator at the second field strength. Here we see the behavior exhibited in some cases in which the quality of $\log G$ fit is good in two places and unsatisfactory in between. This example is a neutron at $\kappa=0.13700$, $m_{\pi}=701$ MeV. Further inspection reveals the red quadrant is associated with a loss of signal in the correlator.